OpenStudy (anonymous):

The difference of a number and 9 is equal to the quotient of 6 and 2

OpenStudy (anonymous):

(n-9) = 6/2 n = 12

OpenStudy (babyslapmafro):

why isn't it (9-n)=6/2

OpenStudy (anonymous):

actually It could. However, the convention is that whichever number is said first is the bigger one. So yes, the mathematical correct way is abs(n-9) = 6/2, which find the other solution